# Context-Aware Coupler Reconfiguration for Tunable Coupler-Based Superconducting Quantum Computers

### v1

@@ 1 Introduction
The performance of NISQ devices heavily relies on the input problem size, specifically the number of qubits in and depth of the quantum circuit. Noise factors, such as thermal relaxation, can hinder quantum computers (QCs) from executing a high quantum circuit depth [9], whereas crosstalk errors reduce parallel gate execution in QCs [10, 11, 12, 13, 14]. Researchers have attempted to mitigate these errors using both hardware and software solutions. At the software level, a compiler can apply problem-specific optimizations [15, 16, 17, 18, 19] tailored to the target hardware. At the hardware level, to address crosstalk errors, tunable connections between qubits, achieved using couplers, can be employed.
When using tunable couplers, the connection between qubit is activated only when a two-qubit gate needs execution; otherwise, it remains inactive. This way, qubits can perform their respective operations without crosstalk errors. However, another type of error is introduced. Ideally, a SQC is assumed to operate as a closed system at absolute zero temperature, an unattainable condition in the real world. Practically, we can only cool the system down to a certain low temperature close to zero, and external interactions are required to control qubits, which introduce errors [20]. Additionally, the tuning of the coupler (on/off) introduces unwanted interactions and errors to the qubits [21].
@@
Figure 1: The NISQ device architectures of IBM’s Q20 Tokyo and Google’s Sycamore are represented where each vertex symbolizes a qubit, and each edge signifies a coupler. Therefore, a controlled-NOT (CNOT) gate can directly execute between any two connected qubits.
The existing methods [22, 23] focus on optimizing the qubit placement layout for application-specific quantum circuits, aiming to optimize connectivity while considering physical constraints, such as frequency-collision effects, to improve fabrication yield rates. However, limited attention has been paid to the configuration of tunable-coupler superconducting quantum computers (SQCs). Lin et al. [24] proposed a method for finding an optimal layout using tunable couplers. However, this method did not consider the physical constraint imposed by frequency-collision effects. Our study bridges the gap by exploring coupling configurations for a specific input quantum circuit. It leverages the contextual connectivity correlation of the quantum circuit to minimize connections between noncorrelated qubits (i.e., without two-qubit gates) while enhancing fidelity and considering to the physical constraints of the SQC. We believe our approach can shed light on the benefits of initializing a layout for tunable coupler SQCs instead of dynamically tuning them during runtime, which can significantly affect qubit reliability.
To achieve this, we must address several key challenges related to the physical constraints of tunable-coupler QCs. Next, we must identify the contextual connectivity patterns within the input quantum circuit to discover the optimal connections. Afterward, we must determine the placement of physical qubits and their connections while ensuring that the final layout complies with the physical constraints of the SQC itself.
@@
2 Background and Motivation
Here, we provide the fundamentals for understanding our proposed method. We discuss the quantum computing basics of qubits, quantum operations, and quantum circuits. Thereafter, we discuss the coupler-based SQC hardware constraints, their physical properties, circuit fidelity, and the motivation behind our proposed method.
(a) Hadamard and CNOT gate
(b) SWAP gate with its equivalent decomposed gates
@@ 2.1 Quantum Computing Basics
Qubits and Quantum Operations: Quantum programs are composed of qubits and gates as the communication operations [25]. Figure 3 shows the quantum basic gates of single and multiple-qubit gates, such as Hadamard (H), Controlled-NOT (CNOT), and SWAP (equivalent to three CNOT) gates. A quantum circuit can be composed of as many qubits as needed denoted as $q$ and represented on horizontal lines in Figure 1(a). The choice of the gates can change the state of the quantum program, and there are two basis states denoted as $|0\rangle$ and $|1\rangle$. Additionally, SWAP gates can be as expensive as three equivalent CNOT gates. We assume that the SWAP gate is considered to have only one weight as the CNOT gate.
Quantum Circuit: An interaction graph (IG) is a graph that represents the interactions between the qubits in a quantum circuit [26, 23, 27]. The nodes in the graph represent the qubits $q$, and the edges in the graph represent the two-qubit gates of $q_{i}$ and $q_{j}$ as $g_{n}$, where $n$ is the number of interactions. To consider a logical circuit, we illustrate the quantum circuit of six qubits ($q_{1}$ to $q_{6}$) with the eight CNOT gates ($g_{1}$ to $g_{8}$) (see Figure 1(a)). For instance, gate $g_{1}$ represents the direct interaction between qubits $q_{1}$ and $q_{2}$, whereas gate $g_{2}$ represents the indirect interaction between qubits $q_{2}$ and $q_{4}$. IGs are employed to identify the qubits that can be placed far apart without affecting the circuit’s performance of the circuit. Consequently, it helps determine the number of qubit connections in the circuit from the logical to the physical level. This information define to the connectivity weight in this study.
@@ 2.2 Coupler-Based SQC Hardware Constraints
Quantum computing hardware can be developed using a variety of technologies, such as SQCs, ion traps, quantum dots, neutral atoms, etc. The SQC is currently one of the most promising technology. Shown in Figure 1 are the architecture of the IBM’s Q20 Tokyo, and Google’s Sycamore. During the NISQ era, all logical qubits must be directly implemented by physical qubits unaided by any QEC codes. This study considers two major limitations and constraints of NISQ hardware as context-aware to the model we employed to design coupler reconfiguration for tunable coupler-based SQCs.
(a) Circuit Analysis
(b) Max Weighted Path Graph Generation
(c) Grid Graph
Generation
(d) Constraint
Configuration
(e) Efficient Coupling Map
Figure 4: Overview of the coupling map generation process. (a) Analysis the contextual correlation between qubits; the coupling weight is assigned depending on the number of correlations between qubits. (b) Generation of the MWPG by connecting the highest weighted path (solid line). (c) and (d) Generation of the grid topology using MWPG and connecting the adjacent and diagonal edges, respectively. (e) Generated qubit physical layout coupling map.
@@
Quantum Circuit Fidelity: Notably, the previous problem is resolved to optimize qubit connection and maximize the overall performance against topologies by IBM and Google. Li et al. [28] reported the average error rates of a single-qubit gate and CNOT gate as $4.43\times 10^{-3}$ and $3.00\times 10^{-2}$, respectively. Due to on-chip placement constraints and routing constraints, couplers can only connect one qubit to its neighboring qubit.  [26, 29, 30]. Besides the physical properties, what we are concerned with the crosstalk error, which affects the quality of circuit fidelity. In this study, we investigate the contextual correlation between qubits, i.e., multi-qubit gates. We do not consider single-qubit gates because they do not incur additional qubit movement overhead during execution on any coupling structure.
@@ 2.3 Motivation
We employ a simple circuit of six qubits and eight gates to explain our motivation for considering qubit mapping, gate error, and different types of qubit architecture for determining the number of physical qubit coupler reconfigurations for tunable coupler-based SQCs.
Quantum circuit’s effect on qubit mapping: When executing a quantum circuit on a QC, the logical qubits need to be mapped to physical qubits on the target quantum architecture. For instance, the constraint of SQC architecture is 2D nearest neighbor qubits. Figure 2 demonstrates a simple example of the quantum circuit with a qubit mapping on different architectures of line and grid graphs.
@@
In this instance, it is clear that a better initial mapping depends on the choice of architecture. Therefore, the bridging gap between the physical qubit graph and the qubit mapping problem can be addressed at the compilation level. This motivates us to propose a compilation method for determining the maximum physical qubit coupler reconfiguration for tunable coupler-based superconducting quantum computers (SQCs). This proposed method can resolve the indirect interactions of the remaining nonexecutable two-qubit gates of the qubit mapping problem, and compensate for the lack of connectivity on the physical qubit connectivity. Thus, this method offers more guaranteed advantages, such as lower compiled circuit depth, SWAP-gate-number reduction and fidelity improvement.
3 Context-Aware Coupler Reconfiguration
@@ 3.1 Overall Architecture
Given that SQC processor adheres to a 2D grid shape [31, 32, 33, 34, 6], we need to reshape the graph from Figure 3(a) to conform to a grid-like structure. This adjustment should also account for the constraints imposed by the SQC’s physical construction, including the transmon qubits and couplers. Each qubit can connect to its neighbors in all directions (grid); however, diagonal edges are prohibited when neighboring connections already include diagonal edges (except diagonal neighbor) [23, 35].
@@ 3.2 Max Weighted Path Graph Generation
       $path\_graph$ = initial empty graph; $sorted\_edge$ = get sort weighted edges in $c\_weight$; for $edge$ in $sorted\_edge$ do
@@
             $n1$, $n2$, $weight$ = $edge$; // max connection validation
@@
As we discussed in the previous section, the topology of the quantum circuit resembles a grid, featuring both adjacent and diagonal edges. Consequently, the configured qubit coupling map topology should follow to this grid-like format. To achieve an efficient grid-like layout, we initiate the process by generating a MWPG using the logical coupling map and correlation data obtained from the analyzed circuit. The MWPG is employed because it can be efficiently transformed into the desired grid-like structure with a specified number of rows and columns. Moreover, it can accommodate qubit layouts characterized by high correlations between individual qubits.
@@ 3.3 Grid Topology Generation
In Figures 3(b) and 3(c), the grid graph is created by arranging nodes adjacent to each other in rows and columns. To initialize the grid graph to construct our proposed MWPG, we select either the first or last node of the path and position them in a row and column format. Thereafter, we connect it to all of the path nodes through edges.
@@ 3.4 Eliminating Constraints on Diagonal Edges
To address the constraints imposed by the physical qubits of the SQC, as discussed in Section 2.2, we must consider which possible configuration of diagonal edges carries greater weight. To achieve this, we partition the diagonal edges into two groups, each accumulating the combined weight of diagonal edges that can coexist within the grid graph. As shown in Figure 6, a grid graph with all possible edges connected is depicted. The diagonal edges are separated into two groups: Group 1 (G1) and Group 2 (G2). G1 contains all the diagonal edges with solid (black) lines, whereas G2 comprises all the diagonal edges with dotted (red) lines. In CA-CORE, such a reconfigured topology graph is only generated when the quantum circuit contains fully connected qubits, meaning that each qubit has correlations with all the other qubits in the quantum circuit.
@@
(a) ‘bigadder_n18’ Circuit
(b) ‘qec9xz_n17’ Circuit
(c) ‘bv_n14’ Circuit
(d) ‘bv_n19’ Circuit
(e) ‘multiplier_n15’ Circuit
(f) ‘multiply_n13’ Circuit
Figure 7: Experiments involve applying a depolarizing noise model to execute the NISQ benchmark quantum circuit from [36] on CA-CORE. These results are then compared with outcomes obtained from various topologies. The collected results are compared with the Exact Simulation’s expectation value, which remains unaffected by noise. Consequently, the fidelity of exact simulation is 1. A result from various topologies closer to 1 indicates a higher fidelity.
4 Evaluation
4.1 Experiment Setup
Benchmarks. We generate random quantum circuits using Qiskit [37], ranging from 10-qubit to 33-qubit configurations. These circuits span up to approximately 200 depths and 3000 gates. This allows us to assess the performance regarding reductions in depth, gates, and SWAP operations. Furthermore, we utilize the low-level QASM benchmarks sourced from [36], which are well-suited for NISQ evaluation and simulation. These benchmarks aid in assessing noise-tolerant performance, given their considerable depth. Moreover, their expectation values can be simulated classically, enabling the extraction of precise results.
Hardware Topology. We compare the compiled CA-CORE with the the existing diversity of fixed-coupling device topologies, such as ‘ibmq_almaden,’ ‘ibmq_cairo,’ and ‘ibmq_prague’ available on Qiskit [37]. Furthermore, we generated a topology inspired by Google’s Sycamore architecture [6].
Algorithm Configuration. Our proposed method was setup to compile circuits to a topology where the number of physical qubits matches the number of logical qubits in the quantum circuit. Couplers are configured only when a correlation (a two-qubit gate) between logical qubits exists. The grid size is adjusted to accommodate based on the required number of qubits.
4.2 Depth, Gate, and SWAP-gate-number Reduction
To determine the number of gates, depth, and SWAP operations for all topologies mentioned in Section 4.1, including CA-CORE, we utilize Qiskit’s transpiler. We configure a custom coupling map setup and set the optimization_level to 00. In this configuration, Qiskit applies minimal optimization, primarily focusing on basic qubit mapping, where logical qubit 0 is mapped to physical qubit 0, and so forth. Subsequently, SWAP gates are inserted for logical qubits that lack connectivity.
Table 1 presents theresults of the performancecomparison between CA-CORE and all benchmark topologies, in terms of depth, gate, and SWAP-gate-number reductions. On average, CA-CORE demonstrates improvements of $13.56\%$, $6.97\%$, and $35.87\%$ for of depth, gate, and SWAP operations, respectively. Specifically, for the 30-qubit circuit, CA-CORE employs the coupling-map configuration depicted in Figure 7(b). In this setup, CA-CORE utilizes diagonal couplings, resulting in reductions of $8.62\%$, $4.46\%$, and $17.30\%$ for depth, gate, and SWAP counts, respectively, in contrast to the result of the Sycamore coupling map, as illustrated in Figure 0(b). The primary distinction between the two configurations lies in the additional diagonal couplings among qubits, demonstrating that integrating these essential coupler connections significantly enhances the performance.

### v2

@@ Abstract
  • March 2024
Keywords: quantum computing, domain-specific application, computer architecture
@@ 1 Introduction
The performance of NISQ devices heavily relies on the input problem size, specifically the number of qubits in and the depth of the quantum circuit. Noise factors, such as thermal relaxation, can hinder quantum computers (QCs) from executing a high quantum circuit depth [9], whereas crosstalk errors reduce parallel gate execution in QCs [10, 11, 12, 13, 14]. Researchers have attempted to mitigate these errors using both hardware and software solutions. At the software level, a compiler can apply problem-specific optimizations [15, 16, 17, 18, 19] tailored to the target hardware. At the hardware level, to address crosstalk errors, tunable connections between qubits, achieved using couplers, can be employed.
When using tunable couplers, the connection between qubits is activated only when a two-qubit gate needs execution; otherwise, it remains inactive. This way, qubits can perform their respective operations without crosstalk errors. However, another type of error is introduced. Ideally, a SQC is assumed to operate as a closed system at absolute zero temperature, an unattainable condition in the real world. Practically, we can only cool the system down to a certain low temperature close to zero, and external interactions are required to control qubits, which introduce errors [20]. Additionally, the tuning of the coupler (on/off) introduces unwanted interactions and errors to the qubits [21].
The existing methods [22, 23] focus on optimizing the qubit placement layout for application-specific quantum circuits, aiming to optimize connectivity while considering physical constraints, such as frequency-collision effects, to improve fabrication yield rates. However, limited attention has been paid to the configuration of tunable-coupler superconducting quantum computers (SQCs). Lin et al. [24] proposed a method for finding an optimal layout using tunable couplers. However, this method did not consider the physical constraint imposed by frequency-collision effects. Our study bridges the gap by exploring coupling configurations for a specific input quantum circuit. It leverages the contextual connectivity correlation of the quantum circuit to minimize connections between noncorrelated qubits (i.e., without two-qubit gates) while enhancing fidelity and considering the physical constraints of the SQC. We believe our approach can shed light on the benefits of initializing a layout for tunable coupler SQCs instead of dynamically tuning them during runtime, which can significantly affect qubit reliability.
To achieve this, we address several key challenges related to the physical constraints of tunable-coupler QCs. Next, we identify the contextual connectivity patterns within the input quantum circuit to discover the optimal connections. Afterward, we determine the placement of physical qubits and their connections while ensuring that the final layout complies with the physical constraints of the SQC itself.
@@
Figure 1: The IBM Q20 Tokyo and Google Sycamore NISQ devices are represented where each vertex symbolizes a qubit, and each edge signifies a coupler. Therefore, a Controlled-NOT (CNOT) gate can directly execute between any two connected qubits.
@@
2 Background and Motivation
Here, we provide the fundamentals for understanding our proposed method. We discuss the basics of quantum computing: qubits, quantum operations, and quantum circuits. Thereafter, we discuss the coupler-based SQC hardware constraints, their physical properties, circuit fidelity, and the motivation behind our proposed method.
@@ 2.1 Quantum Computing Basics
Qubits and Quantum Operations: Quantum programs are composed of qubits and gates as the communication operations [25]. Figure 3 shows the quantum basic gates of single and multiple-qubit gates, such as Hadamard (H), Controlled-NOT (CNOT), and SWAP (equivalent to three CNOT) gates. A quantum circuit can be composed of as many qubits as needed, denoted as $q$, and represented on horizontal lines in Figure 2(a). The choice of the gates can change the state of the quantum program, and there are two basis states denoted as $|0\rangle$ and $|1\rangle$. Additionally, SWAP gates can be as expensive as three equivalent CNOT gates. We assume that the SWAP gate is considered to have only one weight as the CNOT gate.
Quantum Circuit: An interaction graph (IG) is a graph that represents the interactions between the qubits in a quantum circuit [26, 23, 27]. The nodes in the graph represent the qubits $q$, and the edges in the graph represent the two-qubit gates of $q_{i}$ and $q_{j}$ as $g_{n}$, where $n$ is the number of interactions. To consider a logical circuit, we illustrate the quantum circuit of six qubits ($q_{1}$ to $q_{6}$) with the eight CNOT gates ($g_{1}$ to $g_{8}$) (see Figure 2(a)). For instance, gate $g_{1}$ represents the direct interaction between qubits $q_{1}$ and $q_{2}$, whereas gate $g_{2}$ represents the indirect interaction between qubits $q_{2}$ and $q_{4}$. IGs are employed to identify the qubits that can be placed far apart without affecting the circuit’s performance of the circuit. Consequently, it helps determine the number of qubit connections in the circuit from the logical to the physical level. This information defines the connectivity weight in this study.
(a) Hadamard and CNOT gate
(b) SWAP gate with its equivalent
decomposed gates
@@ 2.2 Coupler-Based SQC Hardware Constraints
Quantum computing hardware can be developed using various technologies, such as SQCs, ion traps, quantum dots, neutral atoms, etc. The SQC is currently one of the most promising technology. Shown in Figure 1 are the architecture of the IBM’s Q20 Tokyo, and Google’s Sycamore. During the NISQ era, all logical qubits must be directly implemented by physical qubits unaided by any QEC codes. This study considers two major limitations and constraints of NISQ hardware as context-aware to the model we employed to design coupler reconfiguration for tunable coupler-based SQCs.
@@
Quantum Circuit Fidelity: Notably, the previous problem is resolved to optimize qubit connection and maximize the overall performance against topologies by IBM and Google. Li et al. [28] reported the average error rates of a single-qubit gate and CNOT gate as $4.43\times 10^{-3}$ and $3.00\times 10^{-2}$, respectively. Due to on-chip placement constraints and routing constraints, couplers can only connect one qubit to its neighboring qubit.  [26, 29, 30]. Besides the physical properties, we are concerned with the crosstalk error, which affects the quality of circuit fidelity. In this study, we investigate the contextual correlation between qubits, i.e., multi-qubit gates. We do not consider single-qubit gates because they do not incur additional qubit movement overhead during execution on any coupling structure.
@@ 2.3 Motivation
We employ a circuit of six qubits and eight gates to explain our motivation for considering qubit mapping, gate error, and different types of qubit architecture for determining the number of physical qubit coupler reconfiguration for tunable coupler-based SQCs.
Quantum circuit’s effect on qubit mapping: When executing a quantum circuit on a QC, the logical qubits need to be mapped to physical qubits on the target quantum architecture.
@@
In this case, the choice of architecture significantly impacts the initial mapping. Hence, bridging the gap between the physical qubit graph and the qubit mapping problem is crucial at the compilation level. We propose a compilation method for determining the maximum physical qubit coupler reconfiguration in tunable coupler-based superconducting quantum computers (SQCs). This method addresses indirect interactions of nonexecutable two-qubit gates and compensates for connectivity limitations. It offers advantages such as reduced circuit depth, fewer SWAP gates, and improved fidelity.
3 Proposed Context-Aware Coupler Reconfiguration
@@ 3.1 Overall Architecture
Figure 4: Overview of the coupling map generation process. (a) Analyze the contextual correlation between qubits; the coupling weight is assigned depending on the number of correlations between qubits. (b) Generating the Max Weighted Path Graph (MWPG) by connecting the highest weighted path (solid line). (c) Generating the grid topology using MWPG and connecting adjacent edges. (d) Generating the grid topology using MWPG and connecting diagonal edges. (e) Generated qubit physical layout coupling map.
@@
Given that the SQC processor adheres to a 2D grid shape [6, 31, 32, 33, 34], we need to reshape the graph from Figure 4(a) to conform to a grid-like structure. This adjustment should also account for the constraints imposed by the SQC’s physical construction, including the transmon qubits and couplers. Each qubit can connect to its neighbors in all directions (grid); however, diagonal edges are prohibited when neighboring connections already include diagonal edges (except diagonal neighbor) [23, 35].
@@
       $path\_graph$ = initial empty graph;
      $sorted\_edge$ = get sort weighted edges in $c\_weight$;
      for $edge$ in $sorted\_edge$ do
@@
             $n1$, $n2$, $weight$ = $edge$;
            // max connection validation
@@
3.2 Max Weighted Path Graph Generation
As we discussed in the previous section, the topology of the quantum circuit resembles a grid, featuring both adjacent and diagonal edges. Consequently, the configured qubit coupling map topology should follow this grid-like format. To achieve an efficient grid-like layout, we initiate the process by generating an MWPG using the logical coupling map and correlation data obtained from the analyzed circuit. The MWPG is employed because it can be efficiently transformed into the desired grid-like structure with a specified number of rows and columns. Moreover, it can accommodate qubit layouts characterized by high correlations between individual qubits.
@@ 3.2 Max Weighted Path Graph Generation
3.3 Grid Topology Generation
In Figures 4(b) and 4(c), the grid graph is created by arranging nodes adjacent to each other in rows and columns. To initialize the grid graph to construct our proposed MWPG, we select either the first or last node of the path and position them in a row and column format. Thereafter, we connect it to all of the path nodes through edges.
@@ 3.4 Eliminating Constraints on Diagonal Edges
To address the constraints imposed by the physical qubits of the SQC, as discussed in Section 2.2, we must consider which possible configuration of diagonal edges carries greater weight. To achieve this, we partition the diagonal edges into two groups, each accumulating the combined weight of diagonal edges that can coexist within the grid graph. As shown in Figure 6, a grid graph with all possible edges connected is depicted. The diagonal edges are separated into two groups: Group 1 (G1) and Group 2 (G2). G1 contains all the diagonal edges with solid (black) lines, whereas G2 comprises all the diagonal edges with dotted (red) lines. In CA-CORE, such a reconfigured topology graph is only generated when the quantum circuit contains fully connected qubits, meaning each qubit correlates with all the other qubits in the quantum circuit.
@@
4 Evaluation
4.1 Experiment Setup
Benchmarks. We generate random quantum circuits using Qiskit [36], ranging from 10-qubit to 33-qubit configurations. These circuits span up to approximately 200 depths and 3000 gates. This allows us to assess the performance regarding depth reductions, number of gates, and SWAP operations. Furthermore, we utilize the low-level QASM benchmarks from [37], which are well-suited for NISQ evaluation and simulation. These benchmarks aid in assessing noise-tolerant performance, given their considerable depth. Moreover, their expectation values can be simulated classically, enabling the extraction of precise results.
Hardware Topology. We compare the compiled CA-CORE with the existing diversity of fixed-coupling device topologies, such as ‘ibmq_almaden,’ ‘ibmq_cairo,’ and ‘ibmq_prague’ available on Qiskit [36]. Furthermore, we generated a topology inspired by Google’s Sycamore architecture [6].
Algorithm Configuration. Our proposed method was set up to compile circuits to a topology where the number of physical qubits matches the number of logical qubits in the quantum circuit. Couplers are configured only when a correlation (a two-qubit gate) between logical qubits exists. The grid size is adjusted based on the required number of qubits.
4.1.1 Depth, Gate, and SWAP-gate-number Reduction
To determine the number of gates, depth, and SWAP operations for all topologies mentioned in Section 4.1, including CA-CORE, we utilize Qiskit’s transpiler. We configure a custom coupling map setup and set the optimization_level to 00. In this configuration, Qiskit applies minimal optimization, primarily focusing on trivial qubit mapping, where logical qubit 0 is mapped to physical qubit 0, and so forth. Subsequently, SWAP gates are inserted to establish connectivity for logical qubits that are not directly connected.
@@ 4.1.1 Depth, Gate, and SWAP-gate-number Reduction
Table 1 presents the performance comparison results between CA-CORE and all benchmark topologies regarding depth, gate, and SWAP-gate-number reductions. On average, CA-CORE demonstrates improvements of $13.56\%$, $6.97\%$, and $35.87\%$ for depth, number of gates, and SWAP operations, respectively. Specifically, for the 30-qubit circuit, CA-CORE employs the coupling-map configuration depicted in Figure 8(b). In this setup, CA-CORE utilizes diagonal couplings, resulting in reductions of $8.62\%$, $4.46\%$, and $17.30\%$ for depth, number of gates, and SWAP counts, respectively, in contrast to the result of the Sycamore coupling map, as illustrated in Figure 1(b). The primary distinction between the two configurations lies in the additional diagonal couplings among qubits, demonstrating that integrating these essential coupler connections significantly enhances the performance.
@@
(a) bigadder_n18’ Circuit
(b) qec9xz_n17’ Circuit
(c) bv_n14’ Circuit
(d) bv_n19’ Circuit
(e) multiplier_n15’ Circuit
(f) multiply_n13’ Circuit
Figure 7: Experiments involve applying a depolarizing noise model to execute the NISQ benchmark quantum circuit on CA-CORE. These results are then compared with outcomes obtained from various topologies. The collected results are compared with the Exact Simulation’s expectation value, which remains unaffected by noise. Consequently, the fidelity of exact simulation is 1. A result from various topologies closer to 1 indicates a higher fidelity.